Dakota Fan, Inc., manufactures an inexpensive household fan that it sells to retailers for $25 per unit. All sales are on account, with 30 percent of sales collected in the month of sale and 70 percent collected in the following month. The data that follow were extracted from the companys accounting records. Dakota Fan maintains a minimum cash balance of $20,000. Total payments in January 20x1 are budgeted at $220,000. A schedule of cash collections for January and February of 20x1 revealed the following receipts for the period Cash Receipts January February From December 31 accounts$126,000 From January sales From February sales 87,000 $133,000 5,000 March 20x1 sales are expected to total 8,500 units. Finished-goods inventories are maintained at 30 percent of the following months sales. The December 31, 20x0, balance sheet revealed the following selected figures: cash, $23,600; accounts receivable, $126,000; and finished goods, $24,000. Required: 1. Determine the number of units that Dakota Fan sold in December 20x0 2. Compute the sales revenue for March 20x1 3. Compute the total sales revenue to be reported on Dakota Fans budgeted income statement for the first quarter of 20x1 4. Determine the accounts receivable balance to be reported on the March 31, 20x1, budgeted balance sheet 5. Calculate the number of units in the December 31, 20x0, finished-goods inventory 6. Calculate the number of units of finished goods to be manufactured in January 20x1 7. Calculate the financing required in January, if any, to maintain the firms minimum cash balance.
Answer #1

Solution 1:

Accounts receivables on Dec 31 = $126,000

Total sales of december = $126,000 / 70 % = $180,000

Sales units for december = $180,000 / $25 =7200 units

Solution 2:

Sales units for March 20x1 = Nos of units * Selling price per unit = 8500 * $25 = $212,500

Solution 3:

Sales for January = $87,000 + $133,000 = $220,000

Sales for february = $75,000 / 30% = $250,000

Sales for march = $212,500

Total sales revenue for 1st quarter of 20X1 = $220,000 + $250,000 + $212,500 = $682,500

Solution 4:

Accounts receivables balance to be reported in March 31, 20x1 balance sheet = $212,500 * 70% =$148,750

Solution 5:

Sales units for January = $220,000 / $25 = 8800 units

Finished goods inventory on December 31, 20X0 = 8800 * 30% = 2640 units

Solution 6:

Nos of units of finished goods to be manufacturied in January 20X1 = Sales units for January + Ending inventory for January - Beginning inventory for January

Sales units for february = $250,000 / 25 = 10000 units

Nos of units of finished goods to be manufacturied in January 20X1 = 8800 + (10000*30%) - 2640

= 9160 units

Solution 7:

Cash balance for January before financing = Beginning cash balance + Cash receipt - Cash disbursements

= $23,600 + ($126,000 + $87,000) - $220,000 = $16,600

Minimum cash balance required = $20,000

Required financing = $20,000 - $16,600 = $3,400